OpenStudy (anonymous):

The value of y varies directly with x. What is k the constant of variation for this function, and what is y for the given value of x? y = -1/2 when x = 6. Find y when x = -3

OpenStudy (kropot72):

y = kx k = y/x .............(1) To find the value of k, substitute -1/2 for y and substitute 6 for x in equation (1)

OpenStudy (anonymous):

k = -1/2 right?

OpenStudy (kropot72):

No. You need to simplify the following: \[k=\frac{y}{x}=\frac{-\frac{1}{2}}{6}\]

OpenStudy (anonymous):

woops i meant -1/12

OpenStudy (kropot72):

Correct! Good work. Can you do the second part now?

OpenStudy (kropot72):

\[y=-\frac{x}{12}\]

OpenStudy (anonymous):

youd take y = -1/12(-3) and get 1/4?

OpenStudy (kropot72):

Correct again! Well done :)

OpenStudy (anonymous):

thanks so much (:

OpenStudy (kropot72):

You're welcome :)